OpenStudy (anonymous):

The sum of two consecutive even integers is 226. What is the larger integer? 114 116 112 118

OpenStudy (shinalcantara):

114

OpenStudy (zarkon):

\[\frac{226}{2}+1\]

OpenStudy (anonymous):

how

OpenStudy (zarkon):

the average of the two even numbers will be 226/2 the average is directly between the two even numbers so the smaller number is (226/2)-1 and the larger is (226/2)+1

OpenStudy (shinalcantara):

let x be the smaller integer. since it's consecutive even integer then the larger number would be x+2. then you'll have the equation, x+x+2=226. transposing 2 to the right side and combining the variables you'll have 2x=226-2. you'll get x as 112. since you're looking for the larger number, then you'll have x+2=112+2+114

OpenStudy (shinalcantara):

112+2=114 rather